OpenStudy (anonymous):

find a function y=ax^2+bx+c whose graph has an x intercept of 1, a y intercept of 2, and a tangent line with a slope of -1 at the y intercept

OpenStudy (anonymous):

Let \[ f(x)= a x^2 + b x +c\\ f(0)=2= c\\ f(x)= a x^2 + b x +2 \] We still need to find a and b. Can you do it?

OpenStudy (anonymous):

We have \[ f(1)=0 = a+ b +2\\ a+ b =-2 \] We need another relation between a and b. We still have one property that we have not used.

OpenStudy (anonymous):

\[ f'(x) = 2 a x + b\\ f'(0)= -1= b\\ b=-1\\ \] How much is a?

OpenStudy (anonymous):

a+b =-2 a-1=-2 a=-1 \[ y=f(x)= - x^2 - x +2 \]